OpenStudy (anonymous):

Which expression is equivalent to this one? k · 6 + k · s A. 6(k + s) B. s(6 + k) C. k(6 + s) D. k · (6 + k) · s

OpenStudy (anonymous):

C because you just put the K's together to make 1 and then distibute

OpenStudy (anonymous):

Factoring out k, k(6+s) seems good to me

OpenStudy (anonymous):

hat is the solution to the equation? 7(5 + 9) = 7· x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

I cant see it is a bunch of ???

OpenStudy (anonymous):

The question marks are multiplication *

OpenStudy (anonymous):

7*x+7*9

OpenStudy (anonymous):

so 7(5+9)=7*x+7??

OpenStudy (anonymous):

No its 7(5+9)=7*x+7*9

OpenStudy (anonymous):

ohh okay thanks one second

OpenStudy (anonymous):

Kay

OpenStudy (anonymous):

x=1 did you want me to show you how i got that

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):